# Large Language Models for Designing Participatory Budgeting Rules

### v1

@@ 1. Introduction
There are, however, three challenges to address when employing existing LLM-based methods. (Challenge I) These methods only consider single-objective problems  [^2] By objective, we mean an explicit function inherent to the problem of interest, hence works such as yao2024multi that consider empirical runtime as additional objectives do not apply., whereas we are optimizing for both utility and fairness objectives. (Challenge II) Previous LLM-based methods require a fine-grained performance measure that can capture small improvements during search for ensuring convergence at desirable algorithms romera2024mathematical. In our case, while most utility objectives can provide a rich enough performance measure, since equivalent objectives for quantifying how fair the allocations returned from PB rules are (with respect to some fairness definition) have yet to exist, it is unclear how fairness properties can be effectively enforced for LLM-generated PB rules. (Challenge III) Existing LLM-based methods assume the availability of an efficient evaluation procedure such that the performance of the generated algorithms can be quickly computed. Unfortunately, it is well-documented that verifying whether a given budget allocation satisfies standard fairness properties is intractable (i.e., coNP-complete aziz2017justified; aziz2018complexity). Therefore, even when a suitable fairness objective addressing Challenge II exists, using it to evaluate the “fairness” of LLM-generated PB rules by computing its value for all training problem instances would naturally be intractable as well, rendering the evaluation procedure prohibitively expensive.
@@
In the remainder of this paper, we provide the preliminaries for our considered PB problems in Section 2. Section 3 details our proposed LLMRule framework. We validate its efficacy in Section 4. We discuss related work in Section 5 and conclude our work in Section 6. Readers can also find complete related work in Appendix A, additional theoretical results in Appendix B, complete implementation details in Appendices CF (which include all missing figures and listings), additional experiments involving synthetic data in Appendix G, and limitations and future work in Appendix H.
@@ 2. Preliminaries
We present the classical settings of PB that are most studied in the literature rey2023computational. Let $I=\langle\mathcal{P},c,b\rangle$ be a PB instance, where $\mathcal{P}=\{p_{1},\ldots,p_{m}\}$ is the set of $m$ projects, $c:\mathcal{P}\rightarrow\mathbb{R}_{>0}$ is the cost function that assigns a cost $c(p)\in\mathbb{R}_{>0}$ to every project $p\in\mathcal{P}$, and $b\in\mathbb{R}_{>0}$ is the budget. Note that from a broader perspective, $I$ represents the voting stage of a PB processfootnote 1. For any subset of projects $P\subseteq\mathcal{P}$, we denote $c(P)=\sum_{p\in P}c(p)$ as its total cost.
@@ 3.2. Strong-EJR Approximation
for approval ballots where $S$ is the set of all maximal $(\alpha,P)$-/$P$-cohesive groups $N^{\max}$ from $(I,\bm{A})$ and $s=|S|$, which is bounded by $1,2^{m})$.footnote 4 Clearly, $\phi\in[0,1]$, with 1 being perceived as “fair”. [^5] The upper bound is not necessarily tight since Strong-EJR is unsatisfiable in general aziz2017justified.
@@ 4.1. Experimental Setups
We obtain 617 real-world PB instances from Pabulib [^8] https://pabulib.org/ faliszewski2023pabulib, which were held in the U.S., Canada, Poland, and the Netherlands with either approval or cardinal ballots. We include all instances satisfying the following criteria: having 3 to 25 projects, not “fully funded”footnote 4, and not “experimental” i.e., synthetic. Note that Pabulib only includes a special type of cardinal ballots, called cumulative ballots, in which voters are allocated a specific number of points that they can distribute among the projects ($\sum_{p\in\mathcal{P}}A_{i}(p)\leq 1$). For each of the two types of ballots, we further categorize the dataset into small and large PB instances based on $n$, with small instances having fewer than 1,000 voters. Table 1 specifies our employed dataset. Notably, we train LLMRule using the small instances, then test the learned rules on the large instances, i.e., out-of-distribution (OOD) testing. Due to a plethora of small PB instances with approval ballots, we only use those held in the U.S. for training and the rest for in-distribution (ID) testing.
@@
  • PJR-X[satsat]: MES[satsat] for any DNS function satsatfootnote 10, SeqPhrag, MaximinSupp brill2023proportionality
@@ 4.2. Results
Table 4.2 and Figure 3 show results for the more general setting with cardinal ballots. Surprisingly, LLMRule yields the best utilitarian social welfare among the baselines (beside the utilitarian maximizer), even surpassing the greedy utilitarian rule. Meanwhile, the corresponding Strong-EJR approximation approaches 1, meaning the learned rules are still (empirically) fair. We provide an example of such rule in Figure S7. Compared to the previously shown rule for approval ballots, the greedy scheme from this rule appears to be more sophisticated: after computing the scores, it further ensures projects that bring the highest overall utility are prioritized. Notice that the prompt strategy responsible for producing this rule is newly introduced via our prompt refinement module, which indicates its capability of facilitating the design of novel rules.
@@ 5. Related Work
We discuss the related work focusing on the considered PB settings and LLMs for algorithmic design. For a comprehensive review of the PB literature, we refer readers to the surveys aziz2020participatory; rey2023computational. For LLMs, please refer to liu2024systematic. We follow the notations from Sections 2 and 3 when applicable. Please refer to our complete related work in Appendix A, which expands on the quantification of fairness and existing data-driven approaches for designing PB rules.
@@ Appendix H Limitations and Future Work
We consider PB instances from Pabulibfootnote 8 with at most 25 projects in the experiments, which comprise roughly 67% and 92% of all currently available instances with approval and cardinal ballots, respectively. Despite the significant reduction in complexity for computing Strong-EJR approximation, our techniques described in Section 3.3 may not apply for very large $m$ due to the use of frequent itemset mining algorithms for identifying all maximal cohesive groups. Given that we employ the classical Apriori algorithm huang2000fast, which runs in $O(2^{m})$, future works could consider more advanced algorithms to enhance efficiency and in turn enable LLMRule to work on PB instances with larger $m$. Additionally, leveraging LLMRule to generate PB rules with provable fairness guarantees could be a promising extension.

### v2

@@ 1. Introduction
There are, however, three challenges to address when employing existing LLM-based methods. (Challenge I) These methods only consider single-objective problems [^2] By objective, we mean an explicit function inherent to the problem of interest, hence works (e.g., yao2024multi) that consider empirical runtime as additional objectives do not apply., whereas we are optimizing for both utility and fairness objectives. (Challenge II) Previous LLM-based methods require a fine-grained performance measure that can capture small improvements during search for ensuring convergence at desirable algorithms romera2024mathematical. In our case, while most utility objectives can provide a rich enough performance measure, since equivalent objectives for quantifying how fair the allocations returned from PB rules are (with respect to some fairness definition) have yet to exist, it is unclear how fairness properties can be effectively enforced for LLM-generated PB rules. (Challenge III) Existing LLM-based methods assume the availability of an efficient evaluation procedure such that the performance of the generated algorithms can be quickly computed. Unfortunately, it is well-documented that verifying whether a given budget allocation satisfies standard fairness properties is intractable (i.e., coNP-complete aziz2017justified; aziz2018complexity). Therefore, even when a suitable fairness objective addressing Challenge II exists, using it to evaluate the “fairness” of LLM-generated PB rules by computing its value for all training problem instances would naturally be intractable as well, rendering the evaluation procedure prohibitively expensive.
@@
In the remainder of this paper, we provide the preliminaries for our considered PB problems in Section 2. Section 3 details our proposed LLMRule framework. We validate its efficacy in Section 4. We discuss related work in Section 5 and conclude our work in Section 6. Readers can find complete related work in Appendix A, additional theoretical results in Appendix B, complete implementation details in Appendices CF (which include all missing figures and listings), additional experiments involving synthetic data in Appendix G, and limitations and future work in Appendix H.
@@ 2. Preliminaries
[…]
@@ 3.2. Strong-EJR Approximation
[…]
@@ 4.1. Experimental Setups
[…]
@@
[…]
@@ 4.2. Results
Table 4.2 and Figure 3 show results for the more general setting with cardinal ballots. Surprisingly, LLMRule yields the best utilitarian social welfare among the baselines (beside the utilitarian maximizer), even surpassing the greedy utilitarian rule. Meanwhile, the respective Strong-EJR approximation approaches 1, meaning the learned rules are still (empirically) fair. We provide an example of such rule in Figure S7. Compared to the previously shown rule for approval ballots, the greedy scheme from this rule appears to be more sophisticated: after computing the scores, it further ensures projects that bring the highest overall utility are prioritized. Notice that the prompt strategy responsible for producing this rule is newly introduced via our prompt refinement module, which indicates its capability of facilitating the design of novel rules.
@@ 5. Related Work
We discuss the related work focusing on the considered PB settings and LLMs for algorithmic design. For a comprehensive review of the PB literature, we refer readers to the surveys aziz2020participatory; rey2023computational. For LLMs, please refer to liu2024systematic. We follow the notations from Sections 2 and 3 when applicable. Appendix A expands on the quantification of fairness and existing data-driven approaches for designing PB rules.
@@ Appendix H Limitations and Future Work
[…]